Y=-1/x; (1,-1) (3,-1/3)
Find the equation of the secant line that intersects the given points on the function.

Answer by stanbon(75887)

Y=-1/x; (1,-1) (3,-1/3)
Find the equation of the secant line that intersects the given points on the function.
=====
slope = (-1/3 - (-1))/(3-1) = (2/3)/2 = 1/3
------
Form: y = mx + b
Solve for "b"::
-1 = (1/3)1 + b
b = -4/3
--------
Equation:
y = (1/3)x -(4/3)
-----------------
Cheers,
Stan H.
------------